OpenStudy (sleepyjess):

find an approximate solution to the equation. Round any intermediate work and steps to three decimal places.

OpenStudy (sleepyjess):

\(7ln~(x+2.8)=11.2\)

OpenStudy (mayankdevnani):

Remember :- \[\large \bf \ln(x)=\log_{e} x~~~~where, ~~~e=euler's~constant ≈ 2.71828183\]

OpenStudy (mayankdevnani):

where , e =2.718

OpenStudy (mayankdevnani):

So,suppose :- \[\large \bf \ln(x+2.8)=y\] therefore,we get

OpenStudy (mayankdevnani):

\[\large \bf 7y=11.2\] and find out the value of y first ?

OpenStudy (mayankdevnani):

@sleepyjess

OpenStudy (sleepyjess):

y=1.6 sorry it wouldnt load until now so i couldnt see what you said

OpenStudy (mayankdevnani):

good,so we got, y=1.6

OpenStudy (mayankdevnani):

therefore, \[\large \bf \ln(x+2.8)=1.6\] \[\large \bf \log_{e}(x+2.8)=1.6 \]

OpenStudy (mayankdevnani):

Remember :- \[\huge \bf \log_{x} y=n----->x^n=y\]

OpenStudy (mayankdevnani):

use this property..

OpenStudy (sleepyjess):

\(e^{1.6}=x+2.8\)

OpenStudy (mayankdevnani):

good

OpenStudy (mayankdevnani):

and we have , e=2.7

OpenStudy (mayankdevnani):

plug in the value !!

OpenStudy (mayankdevnani):

what you get ????

OpenStudy (sleepyjess):

4.9=x+2.8 2.1=x

OpenStudy (mayankdevnani):

yes.....

OpenStudy (mayankdevnani):

therefore,x=2.1

OpenStudy (sleepyjess):

@mayankdevnani Sorry to bother you again... I understand everything except where you said 7y=11.2. How did you get this?

OpenStudy (mayankdevnani):

i supposed \[\large \bf \ln(x+2.8)=y\]

OpenStudy (sleepyjess):

Thanks!

OpenStudy (mayankdevnani):

welcome :)